2(a-4)+4=5(9-a)
19-3n=-2n
1/2b-1/2=1/4b
0.6(x-50)=18-0.3(40-10x)

Answer by checkley71(8403)

2(a-4)+4=5(9-a) or 2a-8+4=45-5a or 2a+5a=45+8-4 or 7a=49 or a=49/7 or a=7
19-3n=-2n or 19 = -2n+3n or 19=n.
1/2b-1/2=1/4b or 1/2b-1/4b=1/2 or 1/4b=1/2 or b=1/2 divided by 1/4 or b=2.
0.6(x-50)=18-0.3(40-10x) or 0.6x-30=18-12+3x or 0.6x-3x=18-12+30 or
-2.4x=36 or 2.4x=-36 or x=-15.